National City Corporation, a bank holding company, reported earnings per share (E0) of $2.40 in 2017, and paid dividends per share of $1.06 (D0) in 2017. The earnings were expected to grow 6% (g=6%) annually in the long term. The stock had a beta of 1.05. The risk free rate was 4% and the expected market return was 12%. Estimate the justified leading and trailing P/E Ratios for National City Corporation.
Answer #1

Calculating cost of Equity using CAPM Model,

Cost of Equity = 0.04 + 1.05(0.12 - 0.04)

Cost of Equity = 12.40%

Using Constant Growth Model,

Stock Price = 1.06(1.06)/(0.124 - 0.06)

Stock price = $17.56

Leading PE = 17.56/2.40(1.06)

Leading PE = 6.90

Trailing PE = 17.56/2.4

Trailing PE = 7.32